OpenStudy (anonymous):

Can someone walk me throught this equation? b/3 +1 = 0

OpenStudy (saifoo.khan):

-3

OpenStudy (saifoo.khan):

b = -3

OpenStudy (anonymous):

im going to kill you. i need to learn how to do it.

OpenStudy (saifoo.khan):

Haha.

OpenStudy (saifoo.khan):

ok, look.

OpenStudy (saifoo.khan):

http://www.twiddla.com/574448

OpenStudy (saifoo.khan):

come here.

OpenStudy (anonymous):

b=3*(-1)=-3

OpenStudy (id21):

\[\frac{b}{3}+1=0\] \[\frac{b}{3}=-1\] \[b=(-1)\times3=-3\]

OpenStudy (anonymous):

http://www.twiddla.com/575336

OpenStudy (anonymous):

In this case, a more or less straightforward procedure is to multiply everything by 3 -> b + 3 = 0 -> b = -3

OpenStudy (saifoo.khan):

there?